# Device-Independent Private Quantum Randomness Beacon

### v1

@@ Abstract
The prospects of quantum advantage in secure communication [xu2020secure] and distributed computing [kimble2008quantum] have fueled the implementation of quantum networks [daiss2021quantum, pompili2021realization, chen_integrated_2021, van2022entangling, krutyanskiy2023entanglement, knaut2024entanglement] across the globe in recency. These advantages arise from the users’ ability to detect adversaries and share quantum resources on the network, features that are typically elusive in classical networks. Until now, quantum networks have been used to demonstrate key distribution [chen_integrated_2021], quantum-logic gate operations [daiss2021quantum, pompili2021realization], and entanglement of qubits [van2022entangling, krutyanskiy2023entanglement, knaut2024entanglement]. We introduce a new use-case of quantum networks by showing that quantum networks can be used to generate secure and private random numbers in a cost-efficient manner.
@@
Random numbers are a fundamental resource for modern computing and communication systems. They are vital for numerous essential applications that include cryptography, auditing, simulations, statistical analysis, and even blockchain [amer2025applications]. In particular, a Random Number Generator (RNG) is a critical component of a Hardware Root of Trust, which ensures the security of its cryptographic system’s keys. As such, any open vulnerability of these RNGs may have dire consequences.
@@
Owing to the probabilistic nature of quantum measurements, Quantum RNGs (QRNGs) [ma2016quantum] produce random numbers that are provably secure. In contrast, outputs from RNGs that admit a classical description can be predicted with sufficient information and computation resources. Yet, there exist attacks in which malicious actors gain control over the QRNG output (for a concrete example, we refer the reader to Ref. [smith2021out]). They could do so by changing the quantum-state preparation and measurement processes, thereby invalidating the security proof of the QRNG.
@@
By exploiting the phenomenon of Bell non-locality [bell1964, brunner2014bell], one is able to guarantee the security of the output from a Device-Independent QRNG (DIQRNG) [colbeck2009quantum] without any assumption made on the inner-workings of the devices. Thus, DIQRNG is immune to such attacks by design. However, its implementations [pironio_random_2010, liu2018high, liu_device-independent_2021, shalm2021device] demand apparatus to meet highly stringent benchmarks, requiring DIQRNG users to possess expensive large form-factor devices, which renders DIQRNG impractical.
@@
In this Letter, we identify three desirable properties of QRNGs that are not simultaneously attainable with existing protocols. These desirable properties are (i) that the security of its output can be verified device-independently, (ii) that the QRNG users’ devices can be small form-factor and cost-effective and (iii) that its output must not be correlated with registers possessed by the QRNG provider. The third requirement is especially important in cryptographic applications where the output random numbers are used as the users’ secret keys. This property will ensure that nobody, not even the QRNG provider, will be able to guess the users’ secret keys.
@@
Existing commercial hardware QRNGs would fulfil (ii) and (iii), but not (i). DIQRNG implementations [pironio_random_2010, liu2018high, liu_device-independent_2021, shalm2021device] would fulfill (i) and (iii), but not (ii). The existing commercial QRNG on the cloud [huang2021quantum, kumar2022quantum] would fulfil (ii) but not (i) and (iii). In fact, Ref. [tamura2021quantum] reported that the random numbers generated by the authors on IBM 20Q Poughkeepsie did not pass the statistical test of NIST SP 800-22 [SP800_22], which demonstrates the danger of relying on device characterisation provided by an external vendor. The recent demonstration on certified quantum randomness based on quantum supremacy [liu2025certified] would fulfil (ii) and to some extent, (i), but not (iii). Interestingly, one may also deliver DIQRNG output via cloud but it would still not fulfil (iii). In order to fulfil all three properties, we introduce the Device-Independent Private Quantum Randomness Beacon (DIPQRB).
@@
The DIPQRB utilises the so-called routed Bell test configuration [lobo2024certifying] as illustrated in Fig. 1. Unlike the typical Bell test, we have a router (e.g., an optical switch) that randomly distributes the entangled states between different parties that are involved in the routed Bell test. Furthermore, in the routed Bell test configuration, not all parties need to share non-local correlations to prove quantumness. This will reduce the requirements on the devices of some of the parties.
@@
Crucially, as the Bell test is only performed inside the server, the clients do not need to share non-local correlation with the server. Consequently, the clients do not need detectors that are extremely efficient. Indeed, it was recently showed it is possible to exhibit long-range quantum correlations in a routed Bell test even when the loss in the client’s detector is arbitrarily large [sekatski2025certification]. Most of the cost is concentrated on the server’s side, where a sophisticated setup is required to perform the Bell test. Furthermore, since one server can serve many clients, the cost of the setup required for the Bell test is shared among all the clients in the network.
@@
Our proposed scheme requires a quantum channel to connect the server to each client. This quantum channel can be realised using an optical fibre or a line-of-sight free space quantum link. Remarkably, the scenario that we consider here bears some resemblance to the scenario that is often encountered in quantum key distribution (QKD) [xu2020secure], where an adversary can access the quantum channel to gain some side-information. In this case, it is imperative to analyse the security of our scheme against adversaries that may have quantum side-information. This is different from most of the existing QRNG schemes where the adversary is unable to access the quantum channel. In those cases, there is some flexibility to consider the nature of side-information that is accessible to the adversary.
@@
For example, note that the source, the optical switch, Alice, and Bob are all located inside the server. One can reasonably assume that the server is a secure location which makes it difficult for a potential adversary to physically access. Moreover, the server does not expect to receive any quantum signals from the client. Therefore, one can also use an optical power limiter [zhang2021securing] and an optical isolator to prevent attacks such as the detector blinding attack [lydersen2010hacking] from being launched via the quantum channel. One can then justifiably assume fair sampling, which means that we can safely discard the rounds where Alice or Bob do not produce a conclusive outcome.
@@
One can easily construct a fully device-independent version of the protocol by mapping the inconclusive outcome $\varnothing$ deterministically to the bit value ‘0’. Thus, in the fully device-independent version of the protocol, each party’s output will also be binary, i.e., $\mathcal{A}=\mathcal{B}=\mathcal{C}=\{0,1\}$. As the semi-DI version of the protocol offers a higher randomness generation rate without significantly reducing the level of security since the servers’ devices are inside a secure location, we shall focus on this case. We can consider the case where Alice and Bob perform the Clauser-Horne-Shimony-Holt (CHSH) Bell test [CHSH] inside the server. In the CHSH Bell test, Alice and Bob estimate the so-called winning probability $\omega$, defined as $\omega=\Pr[A_{i}\oplus B_{i}=X_{i}\cdot Y_{i}|S_{i}=0,A_{i}\neq\varnothing,B_{i}\neq\varnothing]$, where $\oplus$ here denotes addition modulo 2.
@@
For the measurement, we assume that each measurement device consists of a pair of threshold single-photon detectors, a polarising beam-splitter and a polarisation modulator which decides the measurement basis. For our simulation, we set Alice’s measurement basis to correspond to a polarisation rotation of $0^{\circ}$ for the setting $x=0$ and $45^{\circ}$ for the setting $x=1$. For Bob’s measurement, the polarisation modulator will rotate the input polarisation by $22.5^{\circ}$ for $y=0$ and $-22.5^{\circ}$ for $y=1$. Lastly, for the client’s device, the polarisation modulator rotates the input polarisation by $0^{\circ}$ for $z=0$ and by $45^{\circ}$ for $z=1$, similar to Alice’s device.
To simplify further, we consider the asymptotic case where the number of rounds are large, i.e., $n\rightarrow\infty$. In this case, we can take the limit $\varepsilon\rightarrow 0$. The randomness generation rate can be approximated by the single-round conditional von Neumann entropy $H(C_{i}|Z_{i},P_{i},E)$. In the practical case with finite number of rounds $n$, there will be a correction term that scales with $O(1/\sqrt{n})$ [metger2024generalised]. This correction term vanishes in the asymptotic limit. More details are given in the Supplementary Material.
@@
We note that the number of measurement bases on the client’s side can be increased to improve the loss tolerance of the protocol [sekatski2025certification]. Remarkably, increasing the number of measurement bases of the client is relatively easy and the cost is relatively low, which further supports the claim that our solution is cost-effective. In particular, this means DIPQRB with more measurement bases for the clients can be implemented with single-photon detectors based on avalanche photodiodes instead of superconducting nanowire single-photon detectors that are significantly more expensive and need to operate at cryogenic temperatures.
@@
Supplementary Material:
Device-Independent Private Quantum Randomness Beacon
In this Supplementary Material, we present the method for entropy estimation for the protocol presented in the main text. As we mentioned in the footnote, in practice, the entropy estimation is typically done by checking if the observed input-output frequency distribution is within some accepted range. We denote the range of accepted distributions as $\mathcal{S}_{\Omega}$. In this Supplementary Material, we want to show that if the observed input-output frequency distribution $\vec{q}$ is within the accepted range, i.e., $\vec{q}\in\mathcal{S}_{\Omega}$, then the conditional smooth min-entropy is higher than some threshold, except with a small probability. In this work, we use the conditional smooth min-entropy as a measure of randomness since it quantifies the amount of private random bits that the client can extract from the protocol.
I Assumptions
@@
The first assumption is necessary since our analysis is based on quantum theory, thus the security of the protocol relies heavily on the correctness of quantum theory. Next, we use the second assumption to greatly simplify the analysis. In recent years, there are new techniques to analyse the security of quantum cryptographic protocols with non-ideal sources of randomness, hence there is a possibility that this assumption can be relaxed. The third assumption is necessary since the client requires the inputs and outputs to the server’s devices to estimate the entropy produced in the protocol. That being said, if the server’s data is leaked after the protocol execution, we claim that the client’s random string will still remain private, thus the server only needs to behave honestly during the protocol execution. The fourth assumption is a standard assumption in routed Bell tests [lobo2024certifying], which form the basis of our protocol. Thus, our protocol naturally inherits this assumption. The fifth assumption is a standard assumption in cryptography. Finally, for the semi-device-independent version of the protocol, we would like to discard rounds where Alice does not register a click, and similarly we would like to discard rounds where the switch directs the photon to Bob and he does not register a click. This is only permissible if whether or not the device registers a click is independent of the input given to that device. We note that this assumption is not necessary in the fully-device-independent framework.
II Generalised entropy accumulation theorem
To estimate the entropy that the protocol generates, we shall use the generalised entropy accumulation theorem (GEAT) with testing [metger2024generalised]. In a nutshell, GEAT is a powerful tool to estimate the entropy produced by a sequence of processes, which obey certain conditions. It can elegantly account for non-independent-and-identical attacks as well as memory effects in the quantum devices. It has been used to analyse a wide-range of quantum cryptographic protocols, especially their finite-size effects. However, the study of the finite-size effects of our protocol is beyond the scope of the current work, and we mainly want to show (using GEAT) that our protocol remains secure even if the adversary launches a sophisticated attack that might depend on the data that she obtains in the preceding rounds, as well as when the devices’ behaviours depend on the previous inputs and outputs. To that end, we want to show that our protocol satisfies the conditions required to apply the GEAT and that the limit of very large number of rounds, the randomness generation rate is lower bounded by the rate that we claim in the main text. We shall leave the finite-key analysis of our protocol for future work.
@@ II Generalised entropy accumulation theorem
    $\mathcal{T}[\omega_{C^{n}E_{n}}]=\sum_{p,q}\left(\Pi_{C^{n}}^{(p)}\otimes\Pi_{E_{n}}^{(q)}\right)\omega_{C^{n}E_{n}}\left(\Pi_{C^{n}}^{(p)}\otimes\Pi_{E_{n}}^{(q)}\right)\otimes{\lvert F(p,q)\rangle\!\langle F(p,q)\rvert}_{D^{n}},$ (2)
@@
The use of the virtual protocol is for the sake of simplicity of the entropy estimation. More concretely, we construct the virtual protocol so that we can directly use the GEAT with testing presented in Ref. [metger2024generalised]. We note that it is possible to avoid this virtual protocol construction by using private test registers (such as the one presented in Ref. [arqand2024generalized]), which will lead to higher certified conditional entropy. This is not our immediate concern since the finite-size effects of the protocol is not within the scope of this Letter, and hence we can set the probability $\gamma$ of choosing a test round to be arbitrarily small.
@@
Now, a few remarks are in order. First, the memory register $R_{i}$ is only associated with the client’s device, Charlie. The server’s devices’ memories are collected by the adversary at the end of each round, via the channel $\mathcal{M}_{i}^{\mathrm{collect}}$. This is to reflect the fact that we want the protocol to be secure even if an adversary has access to the server’s devices. Furthermore, this also helps to satisfy the non-signalling property as the adversary wouldn’t be able to update her side information about $A_{i}$ and $B_{i}$ without having access to the memories of Alice and Bob (which are stored in the quantum registers $Q^{\prime}_{A_{i-1}}$ and $Q^{\prime}_{B_{i-1}}$ and are also collected by $\mathcal{M}_{i-1}^{\mathrm{collect}}$). Next, the construction of the GEAT channel also sheds some light on the level of honesty that we need from the server. The GEAT channel construction requires the following conditions: (1) the preparation of the quantum register $Q_{A_{i}}$ must be independent of the input of the switch $S_{i}$, (2) the preparation of the quantum register $Q_{B_{i}}$ is independent of $X_{i}$ and $Z_{i}$, and similarly, the preparation of the quantum registers $Q_{C_{i}}$ is independent of $X_{i}$ and $Y_{i}$, and lastly, (3) the computation of $D_{i}$ has to be done honestly – in practice, this requires the server to send the registers $A_{i}B_{i}X_{i}Y_{i}S_{i}$ honestly. The first two conditions are the standard non-signalling condition in routed Bell tests [lobo2024certifying]. Therefore, we require the server to behave honestly to respect the routed Bell test non-signalling condition as well as to allow the client to estimate the input-output distribution based on the actual inputs and outputs. Apart from these three conditions, the server’s devices (and the client’s device) may deviate from their expected behaviour.
@@
$\Sigma_{i}(\vec{q})=\left\{\nu_{C_{i}R_{i}D_{i}E_{i}\tilde{E}_{i-1}}=(\mathcal{M}_{i}\otimes\operatorname{id}_{\tilde{E}_{i-1}})[\sigma_{R_{i-1}E_{i-1}\tilde{E}_{i-1}}]:\sigma\in\mathrm{D}(R_{i-1}E_{i-1}\tilde{E}_{i-1}),\,\nu_{D_{i}}=\vec{q}\right\},$ (4)
@@
Informally, the min-tradeoff function (evaluated on the distribution $\vec{q}$) gives a lower bound on the single-round conditional von Neumann entropy evaluated on quantum states whose system $D_{i}$ is distributed according to $\vec{q}$. The technique to construct the min-tradeoff function has been given in Refs. [dupuis2019entropy, metger2024generalised].
Now, let us recall the GEAT with testing derived in Ref. [metger2024generalised]
Theorem 1 (Corollary 4.6 of Ref. [metger2024generalised]).
@@
III Single-round analysis

### v2

@@ I Introduction
The prospects of quantum advantage in secure communication [1] and distributed computing [2] have fueled the implementation of quantum networks [3, 4, 5, 6, 7, 8] across the globe in recency. These advantages arise from the users’ ability to detect adversaries and share quantum resources on the network, features that are typically elusive in classical networks. Until now, quantum networks have been used to demonstrate key distribution [5], quantum-logic gate operations [3, 4], and entanglement of qubits [6, 7, 8]. We introduce a new use-case of quantum networks by showing that quantum networks can be used to generate secure and private random numbers in a cost-efficient manner.
@@
Random numbers are a fundamental resource for modern computing and communication systems. They are vital for numerous essential applications that include cryptography, auditing, simulations, statistical analysis, and even blockchain [9]. In particular, a Random Number Generator (RNG) is a critical component of a Hardware Root of Trust, which ensures the security of its cryptographic system’s keys. As such, any open vulnerability of these RNGs may have dire consequences.
@@
Owing to the probabilistic nature of quantum measurements, Quantum RNGs (QRNGs) [10] produce random numbers that are provably secure. In contrast, outputs from RNGs that admit a classical description can be predicted with sufficient information and computation resources. Yet, there exist attacks in which malicious actors gain control over the QRNG output (for a concrete example, we refer the reader to Ref. [11]). They could do so by changing the quantum-state preparation and measurement processes, thereby invalidating the security proof of the QRNG.
@@
By exploiting the phenomenon of Bell non-locality [12, 13], one is able to guarantee the security of the output from a Device-Independent QRNG (DIQRNG) [14] without any assumption made on the inner-workings of the devices. Thus, DIQRNG is immune to such attacks by design. However, its implementations [15, 16, 17, 18] demand apparatus to meet highly stringent benchmarks, requiring DIQRNG users to possess expensive large form-factor devices, which renders DIQRNG impractical.
@@
In this paper, we identify three desirable properties of QRNGs that are not simultaneously attainable with existing protocols. These desirable properties are (i) that the security of its output can be verified device-independently, (ii) that the QRNG users’ devices can be small form-factor and cost-effective and (iii) that its output must not be correlated with registers possessed by the QRNG provider. The third requirement is especially important in cryptographic applications where the output random numbers are used as the users’ secret keys. This property will ensure that nobody, not even the QRNG provider, will be able to guess the users’ secret keys.
@@
Existing commercial hardware QRNGs would fulfil (ii) and (iii), but not (i). DIQRNG implementations [15, 16, 17, 18] would fulfill (i) and (iii), but not (ii). The existing commercial QRNG on the cloud [19, 20] would fulfil (ii) but not (i) and (iii). In fact, Ref. [21] reported that the random numbers generated by the authors on IBM 20Q Poughkeepsie did not pass the statistical test of NIST SP 800-22 [22], which demonstrates the danger of relying on device characterisation provided by an external vendor. The recent demonstration on certified quantum randomness based on quantum supremacy [23] would fulfil (ii) and to some extent, (i), but not (iii). Interestingly, one may also deliver DIQRNG output via cloud but it would still not fulfil (iii). In order to fulfil all three properties, we introduce the Device-Independent Private Quantum Randomness Beacon (DIPQRB).
@@
The DIPQRB utilises the so-called routed Bell test configuration [24] as illustrated in Fig. 1. Unlike the typical Bell test, we have a router (e.g., an optical switch) that randomly distributes the entangled states between different parties that are involved in the routed Bell test. Furthermore, in the routed Bell test configuration, not all parties need to share non-local correlations to prove quantumness. This will reduce the requirements on the devices of some of the parties.
@@
Crucially, as the Bell test is only performed inside the server, the clients do not need to share non-local correlation with the server. Consequently, the clients do not need detectors that are extremely efficient. Indeed, it was recently showed it is possible to exhibit long-range quantum correlations in a routed Bell test even when the loss in the client’s detector is arbitrarily large [25]. Most of the cost is concentrated on the server’s side, where a sophisticated setup is required to perform the Bell test. Furthermore, since one server can serve many clients, the cost of the setup required for the Bell test is shared among all the clients in the network.
@@
Our proposed scheme requires a quantum channel to connect the server to each client. This quantum channel can be realised using an optical fibre or a line-of-sight free space quantum link. Remarkably, the scenario that we consider here bears some resemblance to the scenario that is often encountered in quantum key distribution (QKD) [1], where an adversary can access the quantum channel to gain some side-information. In this case, it is imperative to analyse the security of our scheme against adversaries that may have quantum side-information. This is different from most of the existing QRNG schemes where the adversary is unable to access the quantum channel. In those cases, there is some flexibility to consider the nature of side-information that is accessible to the adversary.
II Protocol
@@
For example, note that the source, the optical switch, Alice, and Bob are all located inside the server. One can reasonably assume that the server is a secure location which makes it difficult for a potential adversary to physically access. Moreover, the server does not expect to receive any quantum signals from the client. Therefore, one can also use an optical power limiter [26] and an optical isolator to prevent attacks such as the detector blinding attack [27] from being launched via the quantum channel. One can then justifiably assume fair sampling, which means that we can safely discard the rounds where Alice or Bob do not produce a conclusive outcome.
@@ III Example
One can easily construct a fully device-independent version of the protocol by mapping the inconclusive outcome $\varnothing$ deterministically to the bit value ‘0’. Thus, in the fully device-independent version of the protocol, each party’s output will also be binary, i.e., $\mathcal{A}=\mathcal{B}=\mathcal{C}=\{0,1\}$. As the semi-DI version of the protocol offers a higher randomness generation rate without significantly reducing the level of security since the servers’ devices are inside a secure location, we shall focus on this case. We can consider the case where Alice and Bob perform the Clauser-Horne-Shimony-Holt (CHSH) Bell test [28] inside the server. In the CHSH Bell test, Alice and Bob estimate the so-called winning probability $\omega$, defined as $\omega=\Pr[A_{i}\oplus B_{i}=X_{i}\cdot Y_{i}|S_{i}=0,A_{i}\neq\varnothing,B_{i}\neq\varnothing]$, where $\oplus$ here denotes addition modulo 2.
@@
For the measurement, we assume that each measurement device consists of a pair of threshold single-photon detectors, a polarising beam-splitter and a polarisation modulator which decides the measurement basis. For our simulation, we set Alice’s measurement basis to correspond to a polarisation rotation of $0^{\circ}$ for the setting $x=0$ and $45^{\circ}$ for the setting $x=1$. For Bob’s measurement, the polarisation modulator will rotate the input polarisation by $22.5^{\circ}$ for $y=0$ and $22.5^{\circ}$ for $y=1$. Lastly, for the client’s device, the polarisation modulator rotates the input polarisation by $0^{\circ}$ for $z=0$ and by $45^{\circ}$ for $z=1$, similar to Alice’s device.
To simplify further, we consider the asymptotic case where the number of rounds are large, i.e., $n\rightarrow\infty$. In this case, we can take the limit $\varepsilon\rightarrow 0$. The randomness generation rate can be approximated by the single-round conditional von Neumann entropy $H(C_{i}|Z_{i},P_{i},E)$. In the practical case with finite number of rounds $n$, there will be a correction term that scales with $O(1/\sqrt{n})$ [29]. This correction term vanishes in the asymptotic limit. More details are given in the Methods section.
@@
IV Discussion and conclusion
We note that the number of measurement bases on the client’s side can be increased to improve the loss tolerance of the protocol [25]. Remarkably, increasing the number of measurement bases of the client is relatively easy and the cost is relatively low, which further supports the claim that our solution is cost-effective. In particular, this means DIPQRB with more measurement bases for the clients can be implemented with single-photon detectors based on avalanche photodiodes instead of superconducting nanowire single-photon detectors that are significantly more expensive and need to operate at cryogenic temperatures.
@@ IV Discussion and conclusion
V Methods
We now present the method for entropy estimation for the protocol presented in the main text. As we mentioned in the footnote, in practice, the entropy estimation is typically done by checking if the observed input-output frequency distribution is within some accepted range. We denote the range of accepted distributions as $\mathcal{S}_{\Omega}$. In this section, we want to show that if the observed input-output frequency distribution $\vec{q}$ is within the accepted range, i.e., $\vec{q}\in\mathcal{S}_{\Omega}$, then the conditional smooth min-entropy is higher than some threshold, except with a small probability. In this work, we use the conditional smooth min-entropy as a measure of randomness since it quantifies the amount of private random bits that the client can extract from the protocol.
V.1 Assumptions
@@
The first assumption is necessary since our analysis is based on quantum theory, thus the security of the protocol relies heavily on the correctness of quantum theory. Next, we use the second assumption to greatly simplify the analysis. In recent years, there are new techniques to analyse the security of quantum cryptographic protocols with non-ideal sources of randomness, hence there is a possibility that this assumption can be relaxed. The third assumption is necessary since the client requires the inputs and outputs to the server’s devices to estimate the entropy produced in the protocol. That being said, if the server’s data is leaked after the protocol execution, we claim that the client’s random string will still remain private, thus the server only needs to behave honestly during the protocol execution. The fourth assumption is a standard assumption in routed Bell tests [24], which form the basis of our protocol. Thus, our protocol naturally inherits this assumption. The fifth assumption is a standard assumption in cryptography. Finally, for the semi-device-independent version of the protocol, we would like to discard rounds where Alice does not register a click, and similarly we would like to discard rounds where the switch directs the photon to Bob and he does not register a click. This is only permissible if whether or not the device registers a click is independent of the input given to that device. We note that this assumption is not necessary in the fully-device-independent framework.
V.2 Generalised entropy accumulation theorem
To estimate the entropy that the protocol generates, we shall use the generalised entropy accumulation theorem (GEAT) with testing [29]. In a nutshell, GEAT is a powerful tool to estimate the entropy produced by a sequence of processes, which obey certain conditions. It can elegantly account for non-independent-and-identical attacks as well as memory effects in the quantum devices. It has been used to analyse a wide-range of quantum cryptographic protocols, especially their finite-size effects. However, the study of the finite-size effects of our protocol is beyond the scope of the current work, and we mainly want to show (using GEAT) that our protocol remains secure even if the adversary launches a sophisticated attack that might depend on the data that she obtains in the preceding rounds, as well as when the devices’ behaviours depend on the previous inputs and outputs. To that end, we want to show that our protocol satisfies the conditions required to apply the GEAT and that the limit of very large number of rounds, the randomness generation rate is lower bounded by the rate that we claim in the main text. We shall leave the finite-key analysis of our protocol for future work.
@@ V.2 Generalised entropy accumulation theorem
    $\mathcal{T}[\omega_{C^{n}E_{n}}]=\sum_{p,q}\left(\Pi_{C^{n}}^{(p)}\otimes\Pi_{E_{n}}^{(q)}\right)\omega_{C^{n}E_{n}}\left(\Pi_{C^{n}}^{(p)}\otimes\Pi_{E_{n}}^{(q)}\right)\\ \otimes{\lvert F(p,q)\rangle\!\langle F(p,q)\rvert}_{D^{n}},$ (2)
@@
The use of the virtual protocol is for the sake of simplicity of the entropy estimation. More concretely, we construct the virtual protocol so that we can directly use the GEAT with testing presented in Ref. [29]. We note that it is possible to avoid this virtual protocol construction by using private test registers (such as the one presented in Ref. [30]), which will lead to higher certified conditional entropy. This is not our immediate concern since the finite-size effects of the protocol is not within the scope of this Letter, and hence we can set the probability $\gamma$ of choosing a test round to be arbitrarily small.
@@
Now, a few remarks are in order. First, the memory register $R_{i}$ is only associated with the client’s device, Charlie. The server’s devices’ memories are collected by the adversary at the end of each round, via the channel $\mathcal{M}_{i}^{\mathrm{collect}}$. This is to reflect the fact that we want the protocol to be secure even if an adversary has access to the server’s devices. Furthermore, this also helps to satisfy the non-signalling property as the adversary wouldn’t be able to update her side information about $A_{i}$ and $B_{i}$ without having access to the memories of Alice and Bob (which are stored in the quantum registers $Q^{\prime}_{A_{i-1}}$ and $Q^{\prime}_{B_{i-1}}$ and are also collected by $\mathcal{M}_{i-1}^{\mathrm{collect}}$). Next, the construction of the GEAT channel also sheds some light on the level of trust that we need from the server. The GEAT channel construction requires the following conditions: (1) the preparation of the quantum register $Q_{A_{i}}$ must be independent of the input of the switch $S_{i}$, (2) the preparation of the quantum register $Q_{B_{i}}$ is independent of $X_{i}$ and $Z_{i}$, and similarly, the preparation of the quantum registers $Q_{C_{i}}$ is independent of $X_{i}$ and $Y_{i}$, and lastly, (3) the computation of $D_{i}$ has to be done honestly – in practice, this requires the server to send the registers $A_{i}B_{i}X_{i}Y_{i}S_{i}$ honestly. The first two conditions are the standard non-signalling condition in routed Bell tests [24]. Therefore, we require the server to behave honestly to respect the routed Bell test non-signalling condition as well as to allow the client to estimate the input-output distribution based on the actual inputs and outputs. Apart from these three conditions, the server’s devices (and the client’s device) may deviate from their expected behaviour.
@@
$\Sigma_{i}(\vec{q})=\Bigg{\{}\nu_{C_{i}R_{i}D_{i}E_{i}\tilde{E}_{i-1}}=(\mathcal{M}_{i}\otimes\operatorname{id}_{\tilde{E}_{i-1}})[\sigma_{R_{i-1}E_{i-1}\tilde{E}_{i-1}}]:\\ \sigma\in\mathrm{D}(R_{i-1}E_{i-1}\tilde{E}_{i-1}),\,\nu_{D_{i}}=\vec{q}\Bigg{\}},$ (4)
@@
Informally, the min-tradeoff function (evaluated on the distribution $\vec{q}$) gives a lower bound on the single-round conditional von Neumann entropy evaluated on quantum states whose system $D_{i}$ is distributed according to $\vec{q}$. The technique to construct the min-tradeoff function has been given in Refs. [31, 29].
Now, let us recall the GEAT with testing derived in Ref. [29]
Theorem 1 (Corollary 4.6 of Ref. [29]).
@@
V.3 Single-round analysis